A motorboat travels 414km in 6 hours going upstream and 856km in 8 hours going downstream. What is the speed of the boat in still water, and what is the speed of the current?
Answer by stanbon(75887)

A motorboat travels 414km in 6 hours going upstream and 856km in 8 hours going downstream. What is the speed of the boat in still water, and what is the speed of the current?
-------------------------
Let "b" be the speed of the boat and "c" be the speed of the current.
-------------------------
Upstream DATA:
Distance = 414 km ; time = 6 hrs; Rate = d/t = 414/6 = 69 kph
-----------------
Downsteam DATA:
Distance = 856 km ; time = 8 hrs; Rate = d/t = 856/8 = 107 kph
----------------
EQUATION:
b-c = 69
b+c = 107
---------
Add to get:
2b = 176
b = 88 kph (speed of the boat in still water)
----------
Substitute to solve for "c":
88+c = 107
c = 19 kph (speed of the current)
===============
Cheers,
stan H.